# Secure Digital Semantic Communications: Fundamentals, Challenges, and Opportunities

### v2

@@ IV-C Defenses Against Attacks on Deterministic Modulation
Attacks on deterministic modulation are enabled by the explicit quantizer-dequantizer interface and the resulting hard decision boundaries and fixed mapping rules. Defenses should therefore focus on strengthening the discretization interface itself to reduce the chance that small feature perturbations change discrete outcomes, while also improving the robustness of discrete representations as they traverse practical digital physical-layer procedures and packet-based delivery. Practical measures include quantizer and mapping designs with safety margins around decision thresholds, boundary-aware objectives that discourage operating points close to quantization thresholds, and representation formats that reduce the impact of localized errors, for example by using mappings with limited error propagation and adding lightweight error detection for the discrete intermediates. Since deterministic modulation externalizes semantic information into bits, symbols, or indices, integrity protection is also important to prevent direct manipulation of these discrete representations during transmission and packet handling.
For scalar quantization, defenses can emphasize robust bit mapping and protection of the resulting bit stream, for example through Gray-style mappings where appropriate, interleaving or scrambling to spread localized corruption, and error-detecting or error-correcting codes to identify and correct bit flips, insertions, or replays before they translate into large semantic distortions. For symbol quantization, robustness can be improved by enforcing conservative constellation remapping with sufficient separation between decision regions and by ensuring modulation-related configuration is consistently applied across endpoints, so that attackers cannot induce persistent mismatch through configuration manipulation. For vector quantization, protection should include securing codebook distribution and updates, validating codebook versions, and protecting transmitted indices with integrity checks, since small index changes can redirect decoding to a different codeword and cause significant semantic deviation. Finally, because deterministic modulation often uses differentiable approximations during training, systems should be evaluated and refined under inference with hard quantization, so that behavior at deployment under hard quantization is explicitly characterized and strengthened.
@@ IV-D Defenses Against Packet- and Protocol-Level Threats
Packet- and protocol-level defenses in digital SemCom should address two complementary goals: reducing side-channel leakage introduced by packetization and metadata, and preventing active manipulation of packet delivery and protocol state. For privacy, an effective starting point is to minimize and sanitize task-related metadata carried in headers or control signaling and to avoid exposing task identifiers, importance labels, model or codebook versions, and knowledge base synchronization markers in plaintext whenever possible. When such information is required for operation, it should be protected using authenticated encryption and, when feasible, represented in a less revealing form. In addition, resistance to traffic analysis can be improved by shaping observable patterns, for example through padding, batching, or controlled scheduling, so that packet sizes, timing, and retransmission patterns reveal less about task intent or context.
To mitigate active attacks, digital SemCom can rely on standard packet security and protocol hardening measures, while adapting them to the semantic setting where a small number of packets may dominate task performance. Authenticated integrity protection helps prevent packet injection and manipulation, while anti-replay mechanisms and sequence validation mitigate replay and reordering. Robustness against selective dropping can be improved by adding redundancy across packets, using unequal error protection so that task-critical packets receive stronger protection, and employing plausibility checks at the receiver to detect inconsistencies in recovered semantic information when packets are missing or altered. Protocol operations such as acknowledgments, retransmissions, and rate adaptation should be protected against forged feedback by authenticating feedback messages and applying conservative fallback rules when the reported information is inconsistent.
@@ V-C Secure Multiple Access and Scheduling for Discrete Semantic Streams
In digital SemCom, meaning is carried as packet-based bits or symbols, and correct interpretation often depends on shared semantic-side configuration, such as the semantic model version, codebook, or task context agreed between the transmitter and receiver. This makes multiuser operation more tightly coupled to semantic information than in conventional data networks: collisions, imperfect multiuser separation, or mis-scheduling can cause bit or symbol sequences to be attributed to the wrong user or decoded with mismatched configuration, leading to semantic ambiguity even when the physical link remains usable. Adversaries can exploit this coupling by manipulating contention and scheduling to delay or desynchronize configuration updates, injecting packets that imitate another user’s semantic format, or selectively disrupting packets that carry configuration negotiation, so that subsequent packets are processed under mismatched models or codebooks. These risks motivate multiple access and scheduling mechanisms designed for discrete semantic pipelines, including scheduling policies that maintain consistent configuration across users and time, safeguards that reduce cross-user confusion at the bit or symbol stream level, and recovery procedures that restore consistent configuration when anomalies or desynchronization are detected.

### v4

@@ IV-C Defenses Against Attacks on Deterministic Modulation
Attacks on deterministic modulation are enabled by the explicit quantizer-dequantizer interface as well as the resulting hard decision boundaries and fixed mapping rules. Defenses should therefore strengthen the discretization interface to reduce the likelihood that small feature perturbations change discrete outcomes, while improving the robustness of discrete representations as they traverse practical digital physical-layer procedures and packet-based delivery. Practical measures include quantizer and mapping designs with explicit safety margins around decision thresholds, as well as boundary-aware objectives that discourage operating near decision boundaries. They also include representation formats that limit the impact of localized errors, such as mappings with limited error propagation and lightweight error detection for discrete intermediates before semantic decoding.
For scalar quantization, defenses can emphasize robust bit mapping and bitstream protection. Gray-style mappings reduce multi-bit changes caused by a single threshold crossing. Interleaving or scrambling, together with error detection or correction, can mitigate localized corruption before it causes large semantic distortions in downstream semantic decoding. For symbol quantization, robustness can be improved by conservative constellation remapping with sufficient separation between decision regions, and boundary-aware training to reduce sensitivity near decision boundaries. For vector quantization, protection should secure codebook distribution and updates, validate codebook versions, and protect transmitted indices with encryption and integrity checks, since small index changes can redirect decoding to an unintended codeword and cause significant semantic deviation. Codebook integrity monitoring and authenticated updates further reduce the risk of poisoning or unauthorized modifications that manipulate codeword boundaries or endpoint consistency.
@@ IV-D Defenses Against Packet- and Protocol-Level Threats
Packet- and protocol-level defenses in digital SemCom should pursue two complementary goals: reducing side-channel leakage introduced by packetization and metadata, and preventing active manipulation of packet delivery and protocol state. To improve privacy, task-related metadata in headers and control signaling should be minimized and sanitized. In addition, sensitive fields should not be exposed in plaintext whenever possible, such as task identifiers, importance labels, model or codebook versions, and knowledge-base synchronization markers. Resistance to traffic analysis can also be improved by shaping observable patterns through padding, batching, or controlled scheduling. In addition, packetization rules that strongly correlate with semantic importance should be avoided. Otherwise, these correlations may leak packet importance via packet sizes, timing, or retransmission patterns.
To mitigate active attacks, digital SemCom can rely on standard packet security and protocol hardening, while accounting for the semantic setting where a small subset of packets may dominate task performance. Authenticated integrity protection defends against injection and manipulation, and anti-replay mechanisms with sequence validation mitigate replay and reordering for both payload packets and semantic-related control messages. Robustness to selective dropping can be improved through cross-packet redundancy and unequal error protection for task-critical packets. In addition, receiver-side plausibility checks can detect inconsistencies in recovered semantic information under missing or altered packets, particularly when high-impact packets are selectively removed. Fragmentation and reassembly should be validated and hardened. This prevents malformed fragments and ambiguous reassembly outcomes from corrupting semantic information or bypassing sequence validation and integrity checks. Protocol operations such as acknowledgments, retransmissions, and rate adaptation should be protected against forged feedback. This can be achieved by authenticating feedback messages and applying conservative fallback rules when the reported state is inconsistent or statistically implausible. Configuration and synchronization exchanges should be authenticated, including coding and modulation settings, semantic model versions, codebooks, and knowledge base updates. Strict version validation and rollback prevention are also needed. These measures help avoid persistent transmitter-receiver mismatch that can degrade semantic decoding quality over multiple packets.
@@ V-C Secure Multiple Access and Scheduling for Discrete Semantic Streams
In digital SemCom, meaning is carried as packet-based bits or symbols, and correct interpretation often depends on shared semantic-side configuration, such as the semantic model version, codebook, or task context agreed between the transmitter and receiver. This makes multiuser operation more tightly coupled to semantic information than in conventional data networks. Collisions, imperfect multiuser separation, or mis-scheduling can cause bit or symbol sequences to be processed under the wrong semantic context or configuration. This can lead to semantic ambiguity even when user identity is correctly established at the packet level and the physical link remains usable. Adversaries can exploit this coupling in several ways, such as manipulating contention and scheduling to delay or desynchronize configuration updates, injecting packets that imitate another user’s semantic format, or selectively disrupting packets that carry configuration negotiation. As a result, subsequent packets may be processed under mismatched models or codebooks. These risks motivate the design of multiple access and scheduling mechanisms for discrete semantic pipelines. Such mechanisms may include scheduling policies that maintain consistent configuration across users and time, safeguards that reduce cross-user confusion at the bit or symbol stream level, and recovery procedures that restore consistency when anomalies or desynchronization are detected.